# AF Embeddability of the C*-Algebra of a Deaconu-Renault Groupoid

### v1

@@ Abstract.
We study Deaconu-Renault groupoids corresponding to surjective local homeomorphisms on locally compact, Hausdorff, second countable, totally disconnected spaces, and we characterise when the C*-algebras of these groupoids are AF embeddable. Our main result generalises theorems in the literature for graphs and for crossed products of commutative C*-algebras by the integers. We give a condition on the surjective local homeomorphism that characterises the AF embeddability of the C*-algebra of the associated Deaconu-Renault groupoid. In order to prove our main result, we generalise an application of Matui’s HK conjecture by Farsi, Kumjian, Pask, and Sims, that gives an isomorphism from the homology groups of AF algebras and the corresponding K-theory. In this paper, we give an explicit formula for this isomorphism, and we show that this isomorphism is an order isomorphism.
@@ 1. Introduction
A problem that has received much interest during the last decades is the classification of C*-algebras. This problem consists of finding an invariant that identifies all C*-algebras of a certain class up to isomorphism. One of the first results in the area is Glimm’s [22] classification of some uniformly hyperfinite C*-algebras. Glimm’s result was interesting because uniformly hyperfinite algebras are the C*-algebraic version of the $\text{II}_{1}$ factors for von Neumann algebras, but unlike the von Neumann case [42] (see also [12, Theorem V.1.13]), uniformly hyperfinite C*-algebras are not all isomorphic.
Dixmier [15] then generalised this classification by studying a class of C*-algebras called matroid C*-algebras. In [6], Bratteli defined a more general class of C*-algebras, called approximately finite dimensional C*-algebras (or AF algebras). An AF algebra is the completion of an increasing union of finite dimensional C*-algebras. Elliott [19, Theorem 4.3] (see [13, Theorem IV.4.3] for a more modern formulation) classified all AF algebras by their K-theory.
Since Elliott’s classification of AF algebras, different types of C*-algebras have been classified by invariants. Several stably finite C*-algebras [23, 24] are classifiable. Kirchberg [30] and Phillips [43] classified all unital, simple, separable, nuclear and purely infinite C*-algebras (called Kirchberg algebras) satisfying the Universal Coefficient Theorem by their K-theory. See Strung’s book on the classification of C*-algebras [56] for more details.
In certain types of C*-algebras, a dichotomy exists between purely infinite and stably finite. In other words, for certain classes of C*-algebras, a C*-algebra is either stably finite or purely infinite. This dichotomy holds for reduced C*-algebras of certain minimal, topological principal, ample groupoids, Bönicke and Li [5, Theorem D], Rainone and Sims [47, Theorem 7.4], and Ma [37]. Additionally to the classification programme, the possibility of finding a dichotomy for large classes of C*-algebras motivates us to study purely infinite and stably finite C*-algebras, and related properties.
Pimsner and Voiculescu [44] proved that the irrational rotation C*-algebra is embedded into the C*-algebra from [18], which is AF and whose $K_{0}$-group is isomorphic to $\mathbb{Z}+\theta\mathbb{Z}$. This embedding led to the classication of irrational rotation C*-algebras in terms of the angle of rotation. Effros [17] observed the need of investigating subalgebras of AF algebras, since virtually nothing was known about them at the time. Such algebras are denoted to be AF embeddable (or AFE algebras).
Since the formulation of the conjecture that a C*-algebra is AF embeddable if, and only if, it is separable, exact, and quasidiagonal (see [4, Section 7]), the problem of characterising AFE algebras has received much attention. Pimsner [45] characterised the AFE property for crossed products of commutative C*-algebras by the integers. Brown [8] generalised this result to crossed products of AF algebras to the integers, and that the notions of stably finite, quasidiagonal and AF embeddable are equivalent for these crossed products. Brown [7] proved that the crossed product of UHF algebra by $\mathbb{Z}^{k}$ is AF embeddable. Similar results are given in [25, 34, 35, 38]. Liu [36] described the AF embeddability of separable exact quasidiagonal C*-algebras in terms of a condition on the C*-algebras representations. Schafhauser [52, Theorem 6.7] characterised the AF embeddability for topological graph algebras with no sinks, where the notions of stably finite, quasidiagonal and AF embeddable are equivalent. Topological graph algebras are very general and cover the C*-algebras of Deaconu-Renault groupoids, which we study in this paper. So, the equivalence of these notions holds for the C*-algebras we are interested.
In the context of graph algebras, there are several results in the literature studying C*-algebras that are AF, AF embeddable or purely infinite. For example, graph loops characterises when a graph algebra is approximately finite dimensional [32]. Similarly, Schafhauser [54] describes when a graph algebra is AF embeddable, while [3, Proposition 5.3] gives a graph condition that makes the corresponding C*-algebra purely infinite. Szymański [57] showed that every stable, purely infinite, simple and classifiable C*-algebra with $K_{1}$ torsion-free is a graph algebra.
Kumjian and Pask [31] study higher rank graph algebras and give sufficient conditions that make the C*-algebras AF or purely infinite. Clark, an Huef, and Sims [11] characterised the AF embeddability of C*-algebras of row-finite and cofinal $k$-graphs with now sources, for $k=2$. Schafhauser [53, Corollary 6.2] extendded this result for an arbitrary $k$.
In this paper we study C*-algebras of Deaconu-Renault groupoids. These algebras generalise graph algebras [33] and are examples of topological graph algebras [27, Proposition 10.9]. In Subsection 2.1, we describe this class of groupoids in more detail. The main contribution of this paper is the description of the AF embeddability of the C*-algebras of Deaconu-Renault groupoids by a corresponding groupoid property. Our result is analogous to similar results for graph algebras [54], for crossed products of unital commutative C*-algebras by the integers [9, Theorem 11.5], [45, Theorem 9], and for higher rank graphs [11], [53, Corollary 6.2], cited above. In particular, we apply the same strategy from [11], by applying [8, Theorem 0.2] to describe AF embeddability in terms of a K-theory condition, and them analysing this condition to find a more concrete condition which can be described by an equation. We state our main result below.
@@
Note that C*-algebras of Deaconu-Renault groupoids are examples of topological graph algebras. When $X$ is compact, then the corresponding topological graph algebra is unital and, in this case, Theorem 1 is an example of [52, Theorem 6.7], that characterises when a unital topological graph algebra is AF embeddable.
The construction of the Deaconu-Renault groupoid $\mathcal{G}$ depends on the surjective local homeomorphism $\sigma$. The contribution of Theorem 1 is that it provides a condition on $\sigma$ that is equivalent to the AF embeddability of the C*-algebra of the corresponding Deaconu-Renault groupoid.
@@
(Brown) [8, Theorem 0.2] If $B$ is an AF algebra and $\beta\in\mathrm{Aut}(B)$, then the following are equivalent:
@@
Note that AF embeddability, quasidiagonality and the stable finite properties are preserved by stable isomorphisms. If we choose an AF algebra $B$ such that $B\rtimes_{\beta}\mathbb{Z}$ is stably isomorphic to $C^{*}(\mathcal{G})$, then we can apply Brown’s theorem to get a characterisation of the AF embeddability of $C^{*}(\mathcal{G})$. Indeed, we let $c:\mathcal{G}\rightarrow\mathbb{Z}$ be the canonical cocycle, and we let $B=C^{*}(\mathcal{G}(c))$ be the C*-algebra of the skew product $\mathcal{G}(c)$. This C*-algebra is AF by [20, Lemma 6.1]. Theorem III.5.7 of [48] gives the isomorphism
@@
This equivalence gives a characterisation of when the C*-algebra is AF embeddable. Note that item (d) depends on the K-theory, which is very abstract. We use the theory of homology groups of groupoids in order to show that (d) is equivalent to condition (iv) of Theorem 1. By applying results from [20, 39] on homology groups, we get the commutative diagram below.
@@
The existence of the isomorphism $K_{0}(B)\rightarrow H_{0}(\mathcal{G}(c))$ is given by [20, Corollary 5.2], which is an application of Matui’s HK conjecture [39, 40, 41], and is such that $[1_{V}]_{0}\mapsto[1_{V}]$ for every compact open $V\subset\mathcal{G}^{(0)}$. The groupoid $c^{-1}(0)$ is the kernel of the cocycle $c:\mathcal{G}\rightarrow\mathbb{Z}$ defined by $c(x,k,y)=k$. We define the isomorphism $H_{0}(\rho_{\varphi}):H_{0}(\mathcal{G})(c))\rightarrow H_{0}(c^{-1}(0))$ in Section 6. The homomorphisms $K_{0}(\beta)$ and $[\widetilde{\beta}]_{\mathcal{G}(c)}$ are group homomorphisms induced by $\beta$, and the group homomorphism $[\sigma_{\ast}]$ is induced by the map $\sigma_{\ast}$ of Theorem 1.
@@
We need to know the positive elements of $K_{0}(B)$ in order to study item (d). So, when we apply the commutative diagram, we need to guarantee that the isomorphisms above takes positive elements from one group to the another. Although [20, Corollary 5.2] gives the existence of the group isomorphism $K_{0}(B)\cong H_{0}(\mathcal{G}(c))$, we need this isomorphism to be an ordered group isomorphism. We prove that this isomorphism is indeed a ordered group isomorphism using different techniques, and we give an explicit formula for the map, stated as follows:
Definition.
where we identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$ by applying the $j$ map from [48, Proposition II.4.2] and [55, Proposition 9.3.3].
Theorem 2.
  1. 1.
    $\nu([p]_{0})=[\mathrm{tr}_{\varphi_{n}}(p)]$
  2. 2.
  3. 3.
    the isomorphism $\nu$ is precisely the map given by [20, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
This theorem also generalises Matui’s [39, Theorem 4.10], which also gave an order isomorphism for AF groupoids with compact unit spaces, but did not provide a formula for the isomorphism. So, we extend Matui’s result to AF groupoids with non-compact unit spaces. Moreover, by showing that $H_{0}(G)^{+}\cong K_{0}(C^{*}(G))^{+}$, Theorem 2 suggests a stronger version of Matui’s HK conjecture, which considered only group homomorphisms.
We organise the paper as follows: in Section 2, we define Deaconu-Renault groupoids, inductive limits and AF groupoids, presenting some results from the literature. In Section 3, we state the isomorphism by [48, Proposition II.5.1] between the C*-algebra of a skew-product groupoid and its corresponding crossed product. We also define the map $\beta$, used later in the commutative diagram. Then we study the theory of homology groups for groupoids in Section 4. In Section 5, we prove Theorem 2, giving an isomorphism between the homology groups for AF groupoids and the K-theory of the C*-algebras of these groupoids. Similar results have been proved in [20, Corollary 5.2], [39, Theorem 4.10]. Our innovation is that we give an explicit formula for this isomorphism, and we show that it preserves positive elements from one group to another. In Section 6 we apply the results from Sections 4 and 5 to prove the commutative diagram (1.1). In Section 7 we prove Theorem 1, our main result, that describes under certain conditions when the C*-algebra of a Deaconu-Renault groupoid is AF embeddable. Finally, we show in Section 8 that our main theorem generalises some known results in the literature.
@@ 2. Preliminaries
2.1. Deaconu-Renault groupoids
Given the topological spaces $X$ and $Y$, a map $\sigma:X\rightarrow Y$ is a local homeomorphism if every $x\in X$ has an open neighbourhood $\mathcal{U}$ such that $\sigma(\mathcal{U})$ is open and $\sigma|_{\mathcal{U}}:\mathcal{U}\rightarrow\sigma(\mathcal{U})$ is a homeomorphism. A continuous map $\varphi:Y\rightarrow X$ is a continuous section of $\sigma$ if $y=\sigma(\varphi(y))$ for all $y$. Using the argument from the proof of [20, Theorem 4.10], one can show that every surjective local homeomorphism has a continuous section.
Here we use the notation from [58]. It follows from [14, Theorem 1] that the sets $\mathcal{U}_{A,B}^{m,n}$ generate a topology that makes $\mathcal{G}$ a topological groupoids which is locally compact, Hausdorff, second countable, étale. We denote the groupoid $\mathcal{G}$ endowed with this topology a Deaconu-Renault groupoid.
The first example of this groupoid was described in Renault’s book [48, Section III.2] by a class of groupoids whose C*-algebras are Cuntz algebras, where $X$ is the set of sequences $(x_{i})_{i\in\mathbb{N}}$, with $x_{i}\in\{0,1,\dots,n\}$ (or $x_{i}\in\mathbb{N}$ for the Cuntz algebra $\mathcal{O}_{\infty}$), and $\sigma$ was given by $\sigma(x_{0}x_{1}x_{2}\dots)=x_{1}x_{2}\dots$ Deaconu [14] generalises Renault’s definition by considering compact Hausdorff spaces $X$ and assumes that $\sigma:X\rightarrow X$ is a covering map. Under these assumptions, he proved that when $\sigma$ is a homeomorphism, then $C^{*}(\mathcal{G})$ is isomorphic to the crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$. Anantharaman-Delaroche [1, Example 1.2 (c)] generalised it to the case where $\sigma$ is a surjective local homeomorphism, and $X$ is locally compact. She also gave a sufficient conditions to make the C*-algebra of this groupoid purely infinite [1, Propositions 4.2 and 4.3]. Then Renault [49] generalised this definition by considering a map $\sigma$ that is defined only on an open subset of $X$. For the purpose of our paper, here we consider the definition by Anantharaman-Delaroche, and we impose the additional conditions on the space $X$ to be Hausdorff and second countable.
By [49, Proposition 2.9], the groupoid is also amenable. Note that there are different notions of amenability for groupoids: topological amenable and measurewise amenable. By Theorem 3.3.7 of [2], these two notions coincide for groupoids that are locally compact, Hausdorff and étale. In this case, [2, Proposition 6.1.8] implies that the reduced and full C*-algebras of the groupoid coincide.
@@ 2.2. Inductive limits
An inductive sequence in a category $\mathscr{C}$ is a sequence $\{A_{n}\}_{n=1}^{\infty}$ of objects in $\mathscr{C}$ and a sequence $\varphi_{n}:A_{n}\rightarrow A_{n+1}$ of morphisms in $\mathscr{C}$, usually written
@@
Inductive limits, when they exist, are essentially unique in the sense that if $(A,\{\mu_{n}\})$ and $(B,\{\lambda_{n}\})$ are inductive limits of the same inductive sequence, then there is an isomorphism $\lambda:A\rightarrow B$ making the diagram of item (ii) above commutative. See [51, page 93] for details. Because of the essential uniqueness of inductive limits (when they exist), we shall refer to the inductive limit (rather than an inductive limit).
@@ 2.3. AF groupoids
There are different definitions for AF groupoids in the literature. The first was given by Renault [48, Definition III.1.1] in his book, and he proved in [48, Proposition III.1.15] that AF groupoids determine AF algebras up to isomorphism. Because of this result, we can see AF groupoids as the groupoid analogs of AF algebras. Other definitions for such groupoids have been introduced in the literature [50], [21, Definition 3.7], [39, Definition 2.2], and [20, Definition 4.9]. In [10] there is a proof of the equivalence of the definitions [48, Definition III.1.1] and [20, Definition 4.9]. In this paper, we consider the Definition 4.9 of [20], by Farsi, Kumjian, Pask, and Sims.
@@
Note that every AF groupoids is locally compact, Hausdorff, second countable, étale, totally disconnected, and amenable. If $G$ an AF groupoid is the increasing union of the open elementary subgroupoids $G_{n}$. Then $G$ is the inductive limit $G=\displaystyle\lim_{\longrightarrow}G_{n}$ [48, Section III.1].
@@
Given an elementary groupoid $G\cong R(\sigma)$ with $\sigma:X\rightarrow Y$, one we apply the isomorphism to find a local homeomorphism $\widetilde{\sigma}:G^{(0)}\rightarrow Y$ such that $G=R(\widetilde{\sigma})$. So, we can suppose that $G=R(\sigma)$ without loss of generality.
@@
We use the notation $c^{-1}(0)$ because of the following. If $\mathcal{G}$ is the Deaconu-Renault groupoid and $c:\mathcal{G}\rightarrow\mathbb{Z}$ is the canonical cocycle, then $c^{-1}(0)$ is isomorphic to the subgroupoid $\{(x,0,y)\in\mathcal{G}\}$ of $\mathcal{G}$. The isomorphism is defined by $(x,y)\mapsto(x,0,y)$.
3. Skew-products and crossed products
Definition 3.1.
[48, Definition I.1.6] Let $G$ be a groupoid and $c:G\rightarrow\mathbb{Z}$ a homomorphism. Define the sets
respectively. Then $G(c)$ becomes a groupoid, called a skew-product.
Lemma 3.2.
Assume that $G$ is a locally compact, Hausdorff, second countable, étale, measurewise amenable groupoid, and suppose that $c:G\rightarrow\mathbb{Z}$ is a continuous cocycle. We define a dynamical system $\alpha^{c}$ induced by $c$. For every $z\in\mathbb{T}$, $f\in C_{c}(G)$ and $g\in G$, we set $\alpha_{z}^{c}(f)(g)=z^{c(g)}f(g)$. By Proposition II.5.1 of Renault’s book [48], there exists a unique dynamical system defined by the equation above. We make an abuse of notation by denoting this dynamical system $\alpha^{c}:\mathbb{T}\rightarrow\mathrm{Aut}(C^{*}(G))$. The following proposition by Renault [48, Proposition II.5.7] gives an isomorphism between the C*-algebra of skew-product groupoids and the corresponding crossed products. For an introduction to crossed products, see Chapters 1 and 2 of Dana William’s book [59].
Proposition 3.3.
[…]
for $F\in C_{c}(G(c))$, $z\in\mathbb{T}$, $g\in G$.
Now we find a formula for the action $\beta$ on $C^{*}(G(c))$ induced by the dual action $\widehat{\alpha}^{c}$. We will apply this formula in Chapter 7 for Deaconu-Renault groupoids in order to understand in more detail the condition $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$, given by Brown’s theorem.
Proposition 3.4.
Proof.
[…]
$=F(g,a+n).$
By definition, we have that $\rho\rtimes u\circ\widehat{\alpha}^{c}=\beta\circ\rho\rtimes u$. Then $\rho\rtimes u$ is an equivariant isomorphism from $(C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T},\widehat{\alpha}^{c},\mathbb{Z})$ onto $(C^{*}(G(c)),\beta,\mathbb{Z})$. By [59, Lemma 2.65], $C^{*}(G(c))\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic. ∎
4. Homology Groups and K-theory
Our main result, Theorem 1 is an application of Brown’s theorem (Theorem 1.1), which characterises the AF embeddability of a C*-algebra $B\rtimes_{\beta}\mathbb{Z}$ in terms of the K-theory of $B$. Note that condition (iv) is very abstract so, in order to provide a more concrete condition, we need to understand the K-theory of the groupoid C*-algebra $C^{*}(\mathcal{G}(c))$.
The definition of K-theory is very abstract and not very manageable. So, we are going to apply an isomorphism between the K-theory $K_{0}(B)$ and its corresponding homology group $H_{0}(\mathcal{G}(c))$, which allows us to apply several helpful results from the theory of homology groups that will allow us to understand what condition (iv) in Brown’s theorem tells us about the groupoid $\mathcal{G}$.
There is already in the literature an isomorphism between the K-theory and the corresponding homology group for AF algebras, described in [20, Corollary 5.2] by Farsi, Kumjian, Pask and Sims. This isomorphism is not sufficient to study condition (iv) of Brown’s theorem in terms of homology groups because it is not clear if this isomorphism preserves positive elements. So, we are going to extend [20, Corollary 5.2] by showing in Theorem 2 that the isomorphism preserves positive elements, and we give a more explicit formula for the isomorphism. Our idea was to consider the building blocks $R(\sigma)$ of AF groupoids and show that the C*-algebras $C^{*}(R(\sigma))$ are approximated by C*-algebras of the form $C_{0}(V,M_{n})$, whose K-groups are much easier to understand. Theorem 2 also generalises [39, Theorem 4.10] which considers only AF groupoids with compact unit spaces.

### v2

# Positivity of Matui’s HK Conjecture for AF Groupoids

@@ Abstract.
In this paper we generalise an application of Matui’s HK conjecture by Farsi, Kumjian, Pask, and Sims, that gives an isomorphism from the homology groups of AF groupoids to the corresponding K-theory. We give an explicit formula for this isomorphism, and we show that the map is an order isomorphism. Since homology groups are equipped with several useful techniques, this map can help us to understand the K-theory of the C*-algebra in more detail. To illustrate this, we apply the isomorphism to characterise the AF embeddability of the C*-algebra of Deaconu-Renault groupoids.
@@ 1. Introduction
Matui [20] studied how homology groups of groupoids can describe the dynamical properties the topological full groups of a groupoid. There, the homology groups $H_{0}(G)$ and $H_{1}(G)$ for étale groupoids over totally disconnected spaces are described by adapting the construction from Crainic and Moerdijk [8]. In [21, Conjecture 2.6], Matui formulated the HK conjecture, which connects the homology groups with the K-groups of the associated C*-algebras, stated below:
Conjecture 1.1 (Matui’s HK conjecture).
Let $G$ be a locally compact, Hausdorff, étale groupoid, such that $G^{(0)}$ is a Cantor set. Suppose that $G$ is both effective and minimal. Then for $j=0,1$ we have
$K_{j}(C^{*}_{r}(G))\cong\bigoplus_{i=0}^{\infty}H_{2i+j}(G).$
This conjecture holds for AF groupoids with compact unit spaces [20, Theorem 4.10], and for étale groupoids arising from subshifts of finite type [20]. Other examples can be found in [3, 22, 23, 25, 27, 41, 44]. Proietti and Yamashita [28] found an example beyond the original assumptions of the conjecture, while Scarparo [35], Deeley [10], and Ortega and Sanchez [24] provide counterexamples.
Farsi, Kumjian, Pask, and Sims [12] proved Matui’s HK Conjecture for Deaconu-Renault groupoids of ranks one and two, and for k-graph groupoids of single vertex k-graphs which satisfy a mild joint-coprimality condition. Moreover, they generalised Matui’s [20, Theorem 4.10], showing in [12, Corollary 5.2] that the conjecture holds for AF groupoids with non-compact unit spaces.
Additionally, Farsi et al. show that, for an AF groupoid $G$, there exists an isomorphism between the K-group $K_{0}(C^{*}(G))$ and the corresponding homology group $H_{0}(G)$. This map is such that $[1_{V}]_{0}\mapsto[1_{V}]$ for every compact open set $V\subset G^{(0)}$. However, it is not clear from their paper whether this map is an isomorphism of ordered groups, i.e., that it preservers positive maps from one group to the other.
In this paper, we show that the map provided in Farsi et al. for AF groupoids is an ordered group isomorphism. Moreover, we give an explicit formula for the isomorphism, as we state below.
Theorem 1.
  1. 1.
    $\nu([p]_{0})=[\mathrm{tr}_{\varphi_{n}}(p)]$
  2. 2.
  3. 3.
    the isomorphism $\nu$ is precisely the map given by [12, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
The theorem presents two advantages. First, K-groups are important in the theory of C*-algebras, since several results rely on them. However, K-theory is often too abstract and hard to calculate. In contrast, homology groups are more manageable and are equipped with techniques that make them easier to understand. So, for the C*-algebra of an AF groupoid, the explicit formula above can be applied to understand the K-theory in more detail. Second, in some situations, we need to know the positive elements of the $K_{0}$-group, so an application of Matui’s HK conjecture would be useless unless we guarantee that the isomorphism preserves positive elements from one group to another.
Note that Theorem 1 brings another question: for other groupoids where Matui’s HK conjecture hold, is the group isomorphism $K_{0}(C^{*}_{r}(G))\rightarrow\bigoplus_{i=0}^{\infty}H_{2i}(G)$ also an ordered group isomorphism?
In the second part of the paper, we apply Theorem 1 to solve a more practical problem, characterising when the C*-algebra of a large class of Deaconu-Renault groupoids is AF embeddable. The example illustrates the power of homology techniques when one needs to manipulate the elements of the $K_{0}$-group. We state the theorem as follows:
Definition.
where we identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$ by applying the $j$ map from [30, Proposition II.4.2] and [40, Proposition 9.3.3].
@@
The construction of the Deaconu-Renault groupoid $\mathcal{G}$ depends on the surjective local homeomorphism $\sigma$. The contribution of Theorem 2 is that it provides a condition on $\sigma$ that is equivalent to the AF embeddability of the C*-algebra of the corresponding Deaconu-Renault groupoid. This theorem is analogous to similar results for graph algebras [38], for crossed products of unital commutative C*-algebras by the integers [5, Theorem 11.5], [26, Theorem 9]. Schafhauser [39, Theorem C] studied Cuntz-Pimsner algebras. Kumjian and Pask [18] study higher rank graph algebras and give sufficient conditions that make the C*-algebras AF or purely infinite. Clark, an Huef, and Sims [7] characterised the AF embeddability of C*-algebras of row-finite and cofinal $k$-graphs with now sources, for $k=2$. Schafhauser [37, Corollary 6.2] extendded this result for an arbitrary $k$. In particular, we apply the same strategy from [7], by applying [4, Theorem 0.2] to describe AF embeddability in terms of a K-theory condition, and then analysing this condition to find a more concrete condition which can be described by the equation (iv) of the Theorem 2.
Note that C*-algebras of Deaconu-Renault groupoids generalise graph algebras [19] and are examples of topological graph algebras [15, Proposition 10.9]. When $X$ is compact, then the corresponding topological graph algebra is unital and, in this case, Theorem 2 is an example of [36, Theorem 6.7], that characterises when a unital topological graph algebra is AF embeddable.
@@
(Brown) [4, Theorem 0.2] If $B$ is an AF algebra and $\beta\in\mathrm{Aut}(B)$, then the following are equivalent:
@@
Note that AF embeddability, quasidiagonality and stable finiteness are preserved by stable isomorphisms. If we choose an AF algebra $B$ such that $B\rtimes_{\beta}\mathbb{Z}$ is stably isomorphic to $C^{*}(\mathcal{G})$, then we can apply Brown’s theorem to get a characterisation of the AF embeddability of $C^{*}(\mathcal{G})$. Indeed, we let $c:\mathcal{G}\rightarrow\mathbb{Z}$ be the canonical cocycle, and we let $B=C^{*}(\mathcal{G}(c))$ be the C*-algebra of the skew product $\mathcal{G}(c)$. This C*-algebra is AF by [12, Lemma 6.1]. Theorem III.5.7 of [30] gives the isomorphism
@@
This equivalence gives a characterisation of when the C*-algebra is AF embeddable. Note that item (d) depends on the K-theory, which is very abstract. Moreover, in order to understand (d), we need to know the positive elements of $K_{0}(B)$. We use the theory of homology groups of groupoids in order to show that (d) is equivalent to condition (iv) of Theorem 2. By applying results from [12, 20] on homology groups, we get the commutative diagram below.
@@
It follows from [12, Corollary 5.2] that $\mathcal{G}(c)$ is AF groupoid. Then Theorem 1 gives the ordered group isomorphism $K_{0}(B)\rightarrow H_{0}(\mathcal{G}(c))$. The groupoid $c^{-1}(0)$ is the kernel of the cocycle $c:\mathcal{G}\rightarrow\mathbb{Z}$ defined by $c(x,k,y)=k$. We define the isomorphism $H_{0}(\rho_{\varphi}):H_{0}(\mathcal{G})(c))\rightarrow H_{0}(c^{-1}(0))$ in Section 5.3. The homomorphisms $K_{0}(\beta)$ and $[\widetilde{\beta}]_{\mathcal{G}(c)}$ are group homomorphisms induced by $\beta$, and the group homomorphism $[\sigma_{\ast}]$ is induced by the map $\sigma_{\ast}$ of Theorem 2.
@@
We organise the paper as follows: in Section 2, we study inductive limits, AF groupoids, abstract homology groups, and homology groups for groupoids. In Section 3, we study the homology groups of AF groupoids, while in Section 4 we prove Theorem 1, our main result. In Section 5, we apply the isomorphism from Theorem 1 and the tools from homology groups to characterise the AF embeddability of the C*-algebra of the Deaconu-Renault groupoid by Theorem 2. Finally, in Section 6, we show that this theorem generalises known results in the literature.
@@ 2.1. Inductive limits
An inductive sequence in a category $\mathscr{C}$ is a sequence $\{A_{n}\}_{n=1}^{\infty}$ of objects in $\mathscr{C}$ and a sequence $\varphi_{n}:A_{n}\rightarrow A_{n+1}$ of morphisms in $\mathscr{C}$, usually written as
@@
Inductive limits, when they exist, are essentially unique in the sense that if $(A,\{\mu_{n}\})$ and $(B,\{\lambda_{n}\})$ are inductive limits of the same inductive sequence, then there is an isomorphism $\lambda:A\rightarrow B$ making the diagram of item (ii) above commutative. See [34, page 93] for details. Because of the essential uniqueness of inductive limits (when they exist), we shall refer to the inductive limit (rather than an inductive limit).
@@ 2.2. AF groupoids
There are different definitions for AF groupoids in the literature. The first was given by Renault [30, Definition III.1.1] in his book, and he proved in [30, Proposition III.1.15] that AF groupoids determine AF algebras up to isomorphism. Because of this result, we can see AF groupoids as groupoid analogues of AF algebras. Other definitions for such groupoids have been introduced in the literature [32], [13, Definition 3.7], [20, Definition 2.2], and [12, Definition 4.9]. In [6] there is a proof of the equivalence of the definitions [30, Definition III.1.1] and [12, Definition 4.9]. In this paper, we consider the Definition 4.9 of [12], by Farsi, Kumjian, Pask, and Sims.
Given the topological spaces $X$ and $Y$, a map $\sigma:X\rightarrow Y$ is a local homeomorphism if every $x\in X$ has an open neighbourhood $\mathcal{U}$ such that $\sigma(\mathcal{U})$ is open and $\sigma|_{\mathcal{U}}:\mathcal{U}\rightarrow\sigma(\mathcal{U})$ is a homeomorphism. A continuous map $\varphi:Y\rightarrow X$ is a continuous section of $\sigma$ if $y=\sigma(\varphi(y))$ for all $y$. Using the argument from the proof of [12, Theorem 4.10], one can show that every surjective local homeomorphism has a continuous section.
@@
Note that every AF groupoids is locally compact, Hausdorff, second countable, étale, totally disconnected, and amenable. If $G$ an AF groupoid is the increasing union of the open elementary subgroupoids $G_{n}$. Then $G$ is the inductive limit $G=\displaystyle\lim_{\longrightarrow}G_{n}$ [30, Section III.1].
@@
Given an elementary groupoid $G\cong R(\sigma)$ with $\sigma:X\rightarrow Y$, we apply the isomorphism to find a local homeomorphism $\widetilde{\sigma}:G^{(0)}\rightarrow Y$ such that $G=R(\widetilde{\sigma})$. So, we can suppose that $G=R(\sigma)$ without loss of generality.